# StruMamba3D: Exploring Structural Mamba for Self-supervised Point Cloud Representation Learning

### v2

@@ 1 Introduction
Point cloud representation learning aims to extract geometric and semantic features from point clouds to support downstream tasks such as classification, segmentation, and detection. As a foundational task in Computer Vision (CV), it is crucial for various real-world applications, including autonomous driving [19, 4], AR [13], and robotics [24]. However, unlike images or text with structured arrangements, point clouds are inherently unordered and sparse, presenting challenges for designing effective models.
To solve this problem, numerous point cloud representation learning methods have been proposed, which can be broadly divided into three categories according to model architectures: Point-based methods [25, 26, 18, 31, 29, 28], Transformer-based methods [12, 38, 21, 23, 39], and Mamba-based methods [20, 15, 42]. Point-based methods directly process point clouds by modeling the relationships between individual points, such as through set abstraction [25, 26] and graph convolution [18, 31]. These methods offer a simple and computationally efficient design. However, their limited receptive field hinders capturing global information, constraining complex structure modeling. Transformer-based methods leverage attention mechanisms to capture long-range dependencies, expanding the receptive field. Moreover, inspired by self-supervised learning paradigms in NLP [6] and CV [16], many transformer-based methods [38, 21, 23, 39] utilize large unlabeled point cloud datasets for pre-training, achieving impressive results in downstream tasks. However, these methods have limited scalability due to the quadratic complexity of the attention mechanism, as shown in Fig. 1(a). To overcome this limitation, Mamba-based methods [20, 15, 42] first convert the 3D points into a 1D sequence and then utilize selective State Space Model (SSM) [9] with linear complexity to extract point features. Leveraging the robust information aggregation ability of selective SSM, Mamba-based methods have attained promising performance.
@@ 2 Related Work
Self-supervised Representation Learning. To solve the above issues and fully utilize a large number of unlabeled data, self-supervised representation learning methods [36, 43, 38, 23, 21, 39, 3] have emerged. Some contrastive learning-based methods [36, 43] train models to learn distinguishing features by differentiating between positive and negative samples. For instance, PointContrast [36] leverages feature consistency across different views of the same point cloud. Inspired by the success of BERT [6] and MAE [16], many transformer-based methods [38, 23, 21, 39, 3] adopt MPM task for pre-training, encouraging models to infer masked patches from visible ones. PointBERT [38] predicts discrete tokens for masked points, while PointMAE [23] reconstructs masked point clouds. Additionally, recent studies [40, 27] integrate cross-modal information, such as images and text, to enhance pre-training. However, transformer-based methods suffer from quadratic complexity, posing challenges for handling long sequences and making them unsuitable for resource-constrained devices. Recently, SSMs [17, 10, 11, 9, 5] have emerged as a prominent research focus due to their linear complexity and powerful information aggregation ability. In the field of point cloud representation learning, several Mamba-based methods [20, 15, 42] have been proposed. PointMamba [20] employs space-filling curves to serialize point clouds, establishing a simple yet effective baseline. PCM [42] introduces multiple serialization strategies and a bidirectional scanning mechanism to capture comprehensive point cloud structures, but also introduces larger computational overhead. Mamba3D [15] also leverages a bidirectional scanning mechanism and proposes a local norm pooling operation to enhance local geometric features. However, these methods introduce structural distortions when converting 3D points into 1D sequences, limiting model performance, especially in tasks requiring fine-grained geometry like part segmentation. Moreover, they neglect the impact of varying input lengths between pre-training and downstream tasks. Unlike these methods, we propose StruMamba3D, which leverages states as proxies to maintain spatial dependencies among points. Additionally, we introduce a sequence-length adaptive strategy to ensure that the states retain long-sequence memory as the sequence length grows in downstream tasks.
@@ 3.4 Sequence Length-adaptive Strategy
Table 1: Shape Classification on ScanObjectNN and ModelNet40 Datasets. For ScanObjectNN, we report the classification accuracy(%) over the three subsets: OBJ-BG, OBJ-ONLY, and PB-T50-RS. And $\dagger$ indicates using simple rotational augmentation [7] for fine-tuning. For ModelNet40, we report the results with and without the voting strategy.

### v3

@@ 1 Introduction
Point cloud representation learning aims to extract geometric and semantic features from point clouds to support downstream tasks such as classification, segmentation, and detection. As a foundational task in Computer Vision (CV), it is crucial for various real-world applications, including autonomous driving [20, 4], AR [13], and robotics [26, 32]. However, unlike images or text with structured arrangements, point clouds are inherently unordered and sparse, presenting challenges for designing effective models.
To solve this problem, numerous point cloud representation learning methods have been proposed, which can be broadly divided into three categories according to model architectures: Point-based methods [27, 28, 19, 35, 31, 30], Transformer-based methods [12, 42, 22, 25, 44, 43], and Mamba-based methods [21, 15, 47]. Point-based methods directly process point clouds by modeling the relationships between individual points, such as through set abstraction [27, 28] and graph convolution [19, 35]. These methods offer a simple and computationally efficient design. However, their limited receptive field hinders capturing global information, constraining complex structure modeling. Transformer-based methods leverage attention mechanisms to capture long-range dependencies, expanding the receptive field. Moreover, inspired by self-supervised learning paradigms in NLP [6] and CV [16], many transformer-based methods [42, 22, 25, 44] utilize large unlabeled point cloud datasets for pre-training, achieving impressive results in downstream tasks. However, these methods have limited scalability due to the quadratic complexity of the attention mechanism, as shown in Fig. 1(a). To overcome this limitation, Mamba-based methods [21, 15, 47] first convert the 3D points into a 1D sequence and then utilize selective State Space Model (SSM) [9] with linear complexity to extract point features. Leveraging the robust information aggregation ability of selective SSM, Mamba-based methods have attained promising performance.
@@ 2 Related Work
Self-supervised Representation Learning. To solve the above issues and fully utilize a large number of unlabeled data, self-supervised representation learning methods [40, 48, 42, 25, 22, 44, 3] have emerged. Some contrastive learning-based methods [40, 48] train models to learn distinguishing features by differentiating between positive and negative samples. For instance, PointContrast [40] leverages feature consistency across different views of the same point cloud. Inspired by the success of BERT [6] and MAE [16], many transformer-based methods [42, 25, 22, 44, 3] adopt MPM task for pre-training, encouraging models to infer masked patches from visible ones. PointBERT [42] predicts discrete tokens for masked points, while PointMAE [25] reconstructs masked point clouds. Additionally, recent studies [45, 29] integrate cross-modal information, such as images and text, to enhance pre-training. However, transformer-based methods suffer from quadratic complexity, posing challenges for handling long sequences and making them unsuitable for resource-constrained devices. Recently, SSMs [17, 10, 11, 9, 5, 34] have emerged as a prominent research focus due to their linear complexity and powerful information aggregation ability. In the field of point cloud representation learning, several Mamba-based methods [21, 15, 47] have been proposed. PointMamba [21] employs space-filling curves to serialize point clouds, establishing a simple yet effective baseline. PCM [47] introduces multiple serialization strategies and a bidirectional scanning mechanism to capture comprehensive point cloud structures, but also introduces larger computational overhead. Mamba3D [15] also leverages a bidirectional scanning mechanism and proposes a local norm pooling operation to enhance local geometric features. However, these methods introduce structural distortions when converting 3D points into 1D sequences, limiting model performance, especially in tasks requiring fine-grained geometry like part segmentation. Moreover, they neglect the impact of varying input lengths between pre-training and downstream tasks. Unlike these methods, we propose StruMamba3D, which leverages states as proxies to maintain spatial dependencies among points. Additionally, we introduce a sequence-length adaptive strategy to ensure that the states retain long-sequence memory as the sequence length grows in downstream tasks.
@@ 3.4 Sequence Length-adaptive Strategy
Table 1: Shape Classification on ScanObjectNN and ModelNet40 Datasets. For ScanObjectNN, we report the classification accuracy(%) over the three subsets: OBJ-BG, OBJ-ONLY, and PB-T50-RS. And $\dagger$ indicates using simple rotational augmentation [7] for fine-tuning.
@@ 5 Conclusion
6 Acknowledgment
This work is supported by the National Defense Basic Scientific Research program (JCKY2022911B002).
Appendix
7 Overview
In this supplementary material, we first provide additional implementation details (Sec. 8). Next, we present more experimental results (Sec. 9) and qualitative analysis (Sec. 10) to validate and analyze our proposed method. Finally, we discuss the limitations of our approach and potential directions for future research (Sec. 11).
8 More Implementation Details
8.1 Structural SSM Block
In the state-wise update strategy and sequence-length adaptive strategy, we modify the SSM parameter generation process to enhance its efficiency. To provide a more intuitive understanding of the structural SSM block, we present its detailed operations in Algorithm 1. The process begins with the normalization of input points $F_{x}$ and initial spatial states $F_{h}$ followed by their linear projection into $F_{x}^{l}$ and $F_{h}^{l}$. Next, a lightweight convolution is applied to $F_{x}^{l}$ and $F_{h}^{l}$, yielding $\hat{F}_{x}$ and $\hat{F}_{h}$. Before processing $\hat{F}_{x}$ in both forward and backward directions, we compute the relative position offsets $\triangle P$ between the input points $P_{x}$ and spatial states $P_{h}$. Based on the $\hat{F}_{x}$ and $\triangle P$, we generate the SSM parameters ($\mathbf{B}_{o}$, $\mathbf{C}_{o}$, $\mathbf{\Delta}^{\prime}_{o}$). A learnable parameter $\tau$ is introduced to regulate the sampling interval $\mathbf{\Delta}^{\prime}_{o}$, producing $\mathbf{\Delta}_{o}$. Utilizing $\mathbf{\Delta}_{o}$, we transform $\mathbf{\overline{A}}_{o}$, $\mathbf{\overline{B}}_{o}$. Finally, the SSM computes the output point features $\hat{F}_{x}^{o}$ and the updated spatial states $\hat{F_{h}^{o}}$. The forward and backward output point features, $\hat{F}_{x}^{f}$ and $\hat{F}_{x}^{b}$, are gated by $F_{z}$ and then summed together. Similarly, the forward and backward spatial states, $\hat{F}_{h}^{f}$ and $\hat{F}_{h}^{b}$, are aggregated. The final outputs, $F_{x}^{\prime}$ and $F_{h}^{\prime}$, are obtained by connecting the updated inputs and spatial states to $F_{x}$ and $F_{h}$ using residual connections, respectively.
Algorithm 1 Structural SSM Block Process
0: Input point features $F_{x}$: (B, N, D), Initial spatial states $F_{h}$: (B, M, D)
0: Output point features $F_{x}^{\prime}$: (B, N, D), Final spatial states $F_{h}^{\prime}$: (B, M, D)
1:/* normalize the input point features $F_{x}$/
[…]
5:/* process with different direction */
6:for $o\ \text{in}\{\operatorname{forward},\operatorname{backward}\}$ do
7:  /* $\triangle P$ is the relative offsets between the input points and spatial states: (B, N, M, 3) */
8:  $\mathbf{B}_{o}$: (B, N, M) $\leftarrow$ $\operatorname{\phi_{\mathbf{B}}}(\hat{F}_{x})+\operatorname{MLP_{\mathbf{B}}}(\triangle P)$
9:  $\mathbf{C}_{o}$: (B, N, M) $\leftarrow$ $\operatorname{\phi_{\mathbf{C}}}(\hat{F}_{x})+\operatorname{MLP_{\mathbf{C}}}(\triangle P)$
10:  /* softplus ensures positive $\mathbf{\Delta}_{o}$/
11:  $\mathbf{\Delta^{\prime}}_{o}$: (B, N, E) $\leftarrow$ $\log(1+\exp(\phi_{\mathbf{\Delta}}(\hat{F}_{x})))$
12:  /* learnable parameter $\tau$ regulates the total sampling time $\mathbf{\Delta}_{all}$/
13:  $\mathbf{\Delta}_{o}$: (B, N, E) $\leftarrow$ $\tau\times\mathbf{\Delta^{\prime}}_{o}/(\sum_{i=1}^{N}\mathbf{\Delta^{\prime}}_{o}^{i})$
14:  /* $\operatorname{Parameter}^{\mathbf{A}}_{o}$ is learnable parameter: (M, E) */
15:  $\mathbf{\overline{A}}_{o}$: (B, N, M, E) $\leftarrow$ $\mathbf{\Delta}_{o}\otimes\operatorname{Parameter}^{\mathbf{A}}_{o}$
16:  $\mathbf{\overline{B}}_{o}$: (B, N, M, E) $\leftarrow$ $\mathbf{\Delta}_{o}\otimes\mathbf{B}_{o}$
17:  $\hat{F}_{x}^{o}$: (B, N, E), $\hat{F}_{h}^{o}$: (B, M, E) $\leftarrow$ $\mathbf{SSM}(\mathbf{\overline{A}}_{o},\mathbf{\overline{B}}_{o},\mathbf{C}_{o})(\hat{F}_{x},\hat{F}_{h})$
18:end for
19:/* gated linear unit and residual connection */
[…]
23:return  $F_{x}^{\prime}$ and $F_{h}^{\prime}$
8.2 Pre-training Details
Following previous works [25], we pre-train our StruMamba3D on ShapeNet dataset [1], which contains 52,472 unique 3D models across 55 common object categories. For each shape, we sample 1024 points as input and partition them into 64 groups using FPS and KNN, with 60% of the groups being randomly masked. The pre-training process utilizes the AdamW [18] optimizer, incorporating cosine learning rate decay [23]. The initial learning rate is set to 1e-3, with a weight decay of 5e-2 and a dropout rate of 1e-1. The batch size is 128, and training is performed for 300 epochs. The model pre-training is conducted on a single NVIDIA RTX A6000 GPU and takes approximately 18 hours to complete.
8.3 Fine-tuning Details
Shape Classification on ScanObjectNN Dataset. ScanObjectNN is a challenging real-world dataset comprising approximately 15,000 objects across 15 distinct categories. For the shape classification task, we first sample 2048 points from each object as input. Next, we leverage our proposed StruMamba3D to extract discriminative group features, which are then aggregated through maximum and average pooling operations to obtain comprehensive global point cloud representations. These global features are processed through a classification head consisting of three linear layers, with respective output dimensions of 256, 256, and $N_{cls}$ (the number of object categories). Finally, a cross-entropy loss is applied to supervise the predictions.
$\mathcal{L}_{cls}=-\sum_{i=1}^{N_{cls}}y_{i}\log(p_{i}),$ (14)
where $y_{i}$ is the indicator function taking a value of 1 when the true class of the point cloud is $i$ and 0 otherwise, and $p_{i}$ denotes the predicted probability for class $i$. For the fine-tuning process, the model is trained for 300 epochs with a batch size of 32 and a learning rate of 5e-4. We utilize the AdamW optimizer [18] with a weight decay of 5e-2 and apply cosine learning rate decay [23], with an initial learning rate of 1e-6 and a warm-up period of 10 epochs.
Shape Classification on ModelNet40 Dataset. ModelNet40 is a widely used benchmark dataset for 3D shape classification, containing 12,311 unique 3D models spanning 40 categories. For the shape classification task, we begin by sampling 1,024 points from each object as input. All other experimental settings remain consistent with those used for the ScanObjectNN dataset. During testing, we report results both with and without the voting strategy. The voting strategy involves conducting ten tests with random scaling and averaging the predictions to obtain the final result.
Part Segmentation on ShapeNetPart Dataset. ShapeNetPart [1] is a benchmark dataset specifically designed for 3D part segmentation. It comprises 16,881 unique 3D models spanning 16 object categories and 50 part categories, providing a diverse and comprehensive dataset for evaluating fine-grained 3D shape understanding. For the part segmentation task, we begin by sampling 2,048 points from each shape as input. Following PointBERT [42], we adopt a similar segmentation head architecture and extract features from the 444-th, 888-th, and 121212-th layers of our StruMamba3D. These three levels of features are concatenated and then processed separately using average pooling and max pooling to obtain two distinct global features. Besides, leveraging feature propagation from PointNet++ [28], we upsample the concatenated features to match the 2,048 input points, ensuring per-point feature representation. The per-point features are then concatenated with the two global features and passed through a Multi-Layer Perceptron (MLP) for point-wise label prediction. To optimize the model, we use cross-entropy loss to supervise the predictions, which is formulated as follows:
$\mathcal{L}_{seg}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N_{seg}}y^{i}_{j}\log(p^{i}_{j}),$ (15)
where $N$ represents the number of points, $N_{seg}$ denotes the number of part categories, $y^{i}_{j}$ is the indicator function taking a value of 1 when the true class of the point $i$ is $j$ and 0 otherwise, and $p^{i}_{j}$ denotes the predicted probability for point $i$ being classified as $j$. For fine-tuning, the model is trained for 300 epochs with a batch size of 32 and a learning rate of 2e-4. We utilize the AdamW optimizer [18] with a weight decay of 5e-2. Additionally, we apply cosine learning rate decay [23], starting with an initial learning rate of 1e-6 and incorporating a warm-up period of 10 epochs to stabilize early training.
[…]
9 More Experimental Results
9.1 Detailed Results on Part Segmentation
We present per-category part segmentation results on the ShapeNetPart [41] dataset. The results for PointGPT-S [3] and PointMamba [21] are reproduced using their official code. As shown in Tab. 8, our method achieves the best performance in most categories. The segmentation task requires the model to effectively capture the local structure of point clouds. Our approach enhances local structure modeling by introducing spatial states, which play a crucial role in improving segmentation performance. This spatial state modeling allows the network to better preserve spatial relationships and structural details, leading to superior results in segmentation tasks.
Table 8: Detailed Results of Part Segmentation on ShapeNetPart Dataset. We report the mean $\operatorname{IoU}$ for each category. * indicates results reproduced using the official code.
Method $\text{mIoU}_{c}$ $\text{mIoU}_{i}$ airplane bag cap car chair earphone guitar
MaskPoint [22] 84.6 86.0 84.2 85.6 88.1 80.3 91.2 79.5 91.9
PointBERT [42] 84.1 85.6 84.3 84.8 88.0 79.8 91.0 81.7 91.6
PointMAE [42] 84.2 86.1 84.3 85.0 88.3 80.5 91.3 78.5 92.1
PointGPT-S* [3] 84.0 85.9 85.1 86.1 88.8 80.2 91.2 78.8 91.7
PointMamba* [21] 83.8 85.9 84.5 84.7 87.8 80.2 91.2 78.7 91.9
Ours 85.0 86.7 84.8 87.4 89.2 81.3 91.6 80.8 92.2
Method knife lamp laptop motorbike mug pistol rocket skateboard table
MaskPoint [22] 87.8 86.2 95.3 76.9 95.0 85.3 64.4 76.9 81.8
PointBERT [42] 87.9 85.2 95.6 75.6 94.7 84.3 63.4 76.3 81.5
PointMAE [42] 87.4 86.1 96.1 75.2 94.6 84.7 63.5 77.1 82.4
PointGPT-S* [3] 87.0 84.5 95.9 74.2 95.0 83.8 62.6 76.8 81.6
PointMamba* [21] 87.3 85.1 95.8 74.0 94.6 84.7 61.6 75.9 82.4
Ours 88.3 86.0 95.8 78.1 95.5 85.8 63.5 76.9 82.2
9.2 Large-scale 3D Scene Task
To evaluate the scalability of our approach to large-scale 3D scene understanding tasks, we conduct 3D object detection experiments on the ScanNetV2 dataset, which contains approximately 50K points per scene. For fair comparison, we follow the evaluation protocol of Point-M2AE and adopt 3DETR-m as our baseline. Our encoder is configured with the same number of layers as Point-M2AE, processing 2048 input tokens and employing 128 spatial states. We pretrain our model on ScanNetV2 for 1080 epochs using a learning rate of 5e-4 and a batch size of 16. All other hyperparameters are kept consistent with those used in the ShapeNet experiments. During fine-tuning, we strictly follow the 3DETR-m setup to ensure a fair comparison. As shown in Tab. 9, the experimental results clearly demonstrate the effectiveness and generalization ability of our method in large-scale 3D scene settings.
[…]
Method $\text{mAP}_{25}$ $\text{mAP}_{50}$
3DETR 62.1 37.9
3DETR-m 65.0 47.0
Point-M2AE 66.3 48.3
Ours 67.8 50.6
9.3 Ablation of Structural Modeling
Unlike the standard Mamba, we use hidden states to model the structural information of point clouds. During the state update and propagation process, we incorporate the relative positions between input points and states into the generation of SSM parameters. This spatial interaction enables each state to selectively focus on and update the features of points within its designated region. As a result, StruMamba3D eliminates the need for serialization strategies to reorder the point cloud.
To further evaluate the impact of our proposed structural modeling and point serialization strategies, we present the experiments in Tab. 10. For the baseline using the standard Mamba block, adding serialization improves model performance. However, due to the distortion of spatial adjacency, the performance still lags behind our method (89.87 vs. 92.75 on ScanObjectNN, 93.68 vs. 95.06 on ModelNet40, and 84.07 vs. 84.96 on ShapeNetPart). StruMamba3D achieves promising performance without the need for serialization, and adding serialization provides only a modest performance gain. Given the computational overhead introduced by serialization, we choose to remove it from StruMamba3D.
Table 10: Ablation of Structural Modeling.
Backbone Serialization Overall Accuracy mIoUc
ScanNN MN40 SNPart
Mamba 88.24 92.50 82.08
Mamba Z-order 89.28 93.44 83.89
Mamba Hilbert 89.87 93.68 84.07
StruMamba3D 92.75 95.06 84.96
StruMamba3D Z-order 92.47 94.98 84.86
StruMamba3D Hilbert 92.81 94.89 85.12
9.4 Ablation of the Number of Spatial States
In this work, we introduce spatial states into the SSM to capture the structural information of point clouds. To investigate the impact of the number of spatial states on model performance, we conduct ablation studies as shown in Sec. 9.7. The results show that reducing the number of spatial states leads to a performance drop, highlighting the importance of spatial states in capturing point cloud structures. When the number of spatial states exceeds 16, the performance no longer improves significantly with further increases. We believe that 16 spatial states are sufficient to represent the structural information of objects. Therefore, we set $M$=16 as the number of spatial states to balance model performance and computational cost.
Table 11: Ablation of the number $M$ of spatial states.
$M$ Overall Accuracy $\text{mIoU}_{c}$
ScanNN MN40 SNPart
4 91.74 93.35 84.34
8 92.26 94.89 84.71
16 92.75 95.06 84.96
24 92.57 94.81 85.07
9.5 Ablation on the Lightweight Convolution
In the structural SSM block, we propose a lightweight convolution for spatial states and input points. This design offers two key benefits: (1) Enhance feature interactions among spatial states. (2) Replace the causal conv1d in the original SSM block. In lightweight convolution, the number of neighboring points determines the receptive field of the module. To select an appropriate number of neighbors, we conduct an ablation study on the number of neighbors for spatial states and input points. As shown in Tab. 12, the model achieves optimal performance when $k$=4 for spatial states and $k$=8 for input points. Additionally, we observe that the model is more sensitive to the number of neighbors of spatial states than to that of input points. This observation suggests that spatial states play a more crucial role in capturing structural information.
[…]
$k$ for $F_{h}$ $k$ for $F_{x}$ Overall Accuracy $\text{mIoU}_{c}$
ScanNN MN40 SNPart
4 8 92.75 95.06 84.96
2 8 91.91 94.29 84.54
8 8 92.71 94.98 84.92
4 4 92.30 94.65 84.71
4 12 92.37 94.69 84.86
9.6 Linear Complexity of StruMamba3D
Leveraging hardware-aware scan algorithm of Mamba [9], StruMamba3D exhibits linear computational complexity. As illustrated in Fig. 4, we present a FLOPs comparison with the transformer-based method PointMAE, where StruMamba3D demonstrates significantly lower computational cost when processing long sequences.
Figure 4: FLOPs comparison with PointMAE.
We also report the inference time and FPS under different input lengths. As shown in Fig. 5, our method is faster than the Transformer-based method PointGPT, especially with longer input lengths. Besides, our method also achieves competitive inference speed compared to other Mamba-based methods.
Figure 5: Inference time and FPS comparison.
9.7 Performance without Pretraining
We provide the results of training our model from scratch (without pretraining). As shown in following table, our method still performs well, outperforming existing supervised learning only methods.
Method ScanObjectNN ModelNet40 ShapeNetPart
mOA mOA $\text{mIoU}_{c}$
w/o Pretraining 91.33 93.68 83.96
MPM Pretraining 92.09 94.45 84.49
Our Pretraining 92.75 95.06 84.96
10 Qualitative Analysis
10.1 Visualization of Spatial State Correlation
We introduce spatial states into the SSM to capture the local structure of point clouds, thereby preserving the spatial dependencies among points. To validate the effectiveness of spatial states, we visualize the correlations between the output point features and spatial state features produced by the structural SSM. Since neighboring spatial states may focus on the same region, we select five spatial states that are widely spaced for visualization. As shown in Fig. 6, different spatial states focus on different parts of the point cloud. The results demonstrate that our structural SSM can effectively capture the local structure of point clouds by spatial states.
10.2 Visualization of Masked Point Modeling
In the pretraining phase, we adopt masked point modeling as our primary self-supervised learning objective. Given a point cloud, we randomly mask 60% of the point groups and use StruMamba3D to extract features, followed by a shallower decoder that predicts the coordinates of masked points. To further analyze the ability of our model to perceive structural information, we visualize the reconstruction results on the ShapeNet dataset. As shown in Fig. 7, our model can effectively reconstruct the original structure even when 60% of the points are masked. Moreover, our model performs remarkably in reconstructing complex geometric patterns, such as the delicate shapes of chair backs and table legs. These visualization results validate that our approach can effectively capture and encode both local and global structural information from point clouds, which is crucial for downstream tasks.
10.3 Visualization of Part Segmentation
In this subsection, we present the qualitative results of part segmentation on the ShapeNetPart validation set, comparing ground truth and predictions. As shown in Fig. 8, our method demonstrates highly competitive performance in part segmentation. Notably, the ShapeNetPart dataset contains certain annotation errors, as highlighted by the red circles. Nevertheless, our method not only achieves a high mIoU but also correctly segments points with erroneous annotations. Moreover, for complex structures such as wheels and chair legs, our approach delivers accurate and reliable segmentation results.
11 Discussions
In this section, we discuss the limitations of our work and potential directions for future research. Our primary goal is to fully exploit the potential of Mamba for point cloud representation learning. To address two key issues of Mamba: disrupting the spatial adjacency of point clouds and struggling to maintain long-sequence memory in downstream tasks, we propose the structural SSM block and the sequence length-adaptive strategy, respectively. While our approach effectively models the structural information of point clouds and achieves significant performance improvements across four downstream tasks, certain challenges remain. Specifically, in part segmentation tasks, we observe some results with unclear boundaries. We attribute this limitation to the nature of single-scale models, which focus on fixed-scale features and struggle to capture geometric details across multiple scales. For example, larger structures such as chair backs and finer details like table legs require different scale features for precise differentiation. To address this, a promising future direction is the development of a hierarchical Mamba that can perform multi-scale point cloud feature extraction through spatial state modeling. This could enhance the ability of the model to adaptively capture both global structures and fine-grained details, leading to more accurate segmentation results.
Figure 6: Visualization of correlation between the output point and spatial state features produced by the structural SSM. For each point cloud, we select five spatial states that are widely spaced for visualization.
Figure 7: Visualization of reconstructed masked regions on ShapeNet. Full points represent the raw point cloud, while visible points correspond to the input points with 60% masked. The reconstruction results consist of the predicted points combined with visible points.
Figure 8: Visualization of part segmentation of our Structural Mamba on ShapeNetPart. Different colors represent different parts. The red circles highlight points with obvious annotation errors.